using the property of exponents can you write the complex number in standard form please =]
3+i^3

Answer by stanbon(48545)

using the property of exponents can you write the complex number in standard form please =]
3+i^3
--------------
i^3 = (i)^2*i = -i
-------------------
3+i^3 = 3-i
===================
Cheers,
Stan H.